The area of a triangle must be at most 24 inches, the base is 6 inches, and the height is x inches. Find the possible values for x.
A. 0 B. x<8
C. 0 D. 0

Answer by venugopalramana(3286)

AREA = (1/2)*BASE*HEIGHT…..AREA=24 AT MOST
HEIGHT=2*AREA/BASE=2*24/6=8 ATMOST=X
0